In 7 years from now a company plans replacing machinery that now costs 240,000. The company wants to save for its replacement beginning now by depositing 10,000 now to open an account followed by 7 equal amounts at the end of each year with the last deposit made when the replacement machine is purchased. The account has an interest rate of 5.50%. The expected annual price increase of the machines is 2.85%.

a. What is the forecasted cost of the machine when it is replaced applying the price increase rate?

b. What annual deposit amount is required to have the funds required available for machines purchase?
Answer #1

a

Future value = present value*(1+ rate)^time
Future value = 240000*(1+0.0285)^7
Future value = 292173.83

b

Future value = present value*(1+ rate)^time
Future value = 10000*(1+0.055)^7
Future value = 14546.79

Amount financed by equal payments = 292173.83-14546.79=277627.04

FVOrdinary Annuity = C*(((1 + i )^n -1)/i)
C = Cash flow per period
i = interest rate
n = number of payments
277627.04= Cash Flow*(((1+ 5.5/100)^7-1)/(5.5/100))
Cash Flow = 33582.99 = annual deposit